OpenStudy (anonymous):

Solve each system by graphing y=3/2x-9 y=3x+9 and if you show how it looks when you graph it too that would be nice

sam (.sam.):

Where you got stucked?

OpenStudy (anonymous):

I just need someone to explain the whole problem for me you know the whole thing in numbers step by step

sam (.sam.):

You can plot these equations by substituting: When x=-5,-4,-3,-2,-1,0,1,2,3,4,5 y=?,?,?,?,?,?,?,?,?,?

sam (.sam.):

There are 2 graphs you'll have to plot y=3/2x-9 and y=3x+9

OpenStudy (anonymous):

I mean just write all the numbers down step by step.

sam (.sam.):

Once you got it, it looks like

OpenStudy (anonymous):

Ok how about you solve half of the problem than explain the rest to me.

OpenStudy (anonymous):

You should check out mathway.com it works out these problems :)

sam (.sam.):

You have to find some coordinates to plot so we'll use some values for y y=3/2x-9 When y=-5 x=8/3 Then we have coordinate (8/3,-5) Find more coordinates to plot